OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (anonymous):

y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

Use the slope formula

OpenStudy (anonymous):

y − 2 = −one fifth(x + 3) is the required answer.

OpenStudy (anonymous):

Thank you, how though?

OpenStudy (anonymous):



OpenStudy (anonymous):

we can find the slope fromabove points that is -1/5 and use formula y=mx+c

OpenStudy (anonymous):